# NSNQuant: A Double Normalization Approach for Calibration-Free Low-Bit Vector Quantization of KV Cache

### v1

@@ 4.2 PPL evaluation
Table 2: Perplexity on WikiText-2 and C4 with a context length of 4096. The results of CQ reported in the original paper are marked with $\dagger$.
@@
Table 2 presents the perplexity (PPL) of various models evaluated on WikiText2 and C4. Although our framework uses full-precision key and value during the prefilling stage, we follow the evaluation protocol of CQ and KVQuant by quantizing them all in this experiment. Despite its difference from the generation scenario, we find PPL evaluation useful because it measures quantization quality for a sequence with a single forward pass. Furthermore, the results are aligned with those in Table 15, where we report the PPL evaluation results in the generation scenario with residuals. Therefore, we report PPL results in the main table, and use them in ablation studies.
@@ 4.3 LongBench evaluation
Table 3: Evaluation results on LongBench. The task subset is selected following KIVI (liu2024kivi, ). More results with different models can be found in Table 16.
Model Method Bits Qasper QMSum MultiNews TREC TriviaQA SAMSum LCC RepoBench-P Avg.
@@ 4.4 Evaluation on more datasets
4.5 Efficiency analysis
To evaluate time and memory efficiency of NSNQuant, we measure memory usage and throughput of NSNQuant. Following previous works (liu2024kivi, ; kwon2023efficient, ), we use the synthetic data to simulate the scenario of ShareGPT (sharegpt2023, ) where the average input length is 161 and the average generation length is 338. The evaluation is performed on a linux server with a single A100-80GB GPU with LLaMA2-7B.
The result is illustrated in Figure 4. While FP16 baseline suffers from OOM for large batch sizes, NSNQuant-2b and NSNQuant-1b scale efficiently to larger batch sizes, achieving 4×4$\times$4 × larger batch sizes and 3×3$\times$3 × speedup in throughput. Note that the memory increases are not proportional to average bits since the residual is also included in the memory usage.
@@ 4.6 Ablation study
The Figure 5 shows the cosine similarity between the original vector $v$ and the quantized vector $v_{Q}:=\mathbb{C}[\operatorname{VQ}(v)]$ with different codebooks. The fine-tuned version gives higher cosine similarities compared to the K-Means codebook. Furthermore, the cosine similarities are consistently high in the early layers, although the standardization of NSN does not hold in these layers. Notably, the cosine similarities measured using the key-value tokens (lines with markers) are very similar to those measured using synthetic standard normal data (dotted lines). This suggests that NSN successfully aligns the output distribution with the standard normal distribution, and the codebook trained only on synthetic data effectively quantizes such outcomes. As a result, codebook tuning improves perplexity from 5.294 to 5.285 with NSNQuant-2b and 6.910 to 6.703 with NSNQuant-1b.
@@ 5 Conclusion
In this work, we propose NSNQuant, a calibration-free vector quantization (VQ) method for compressing KV cache of LLMs. NSNQuant effectively aligns the token distribution with the standard normal distribution through the NSN (Normalize-Shift-Normalize) process, enabling the use of the specialized codebook for the standard normal distribution. Through extensive experiments, we demonstrate that unlike calibration-based VQ methods, NSNQuant generalizes well across different tasks and datasets. In particular, NSNQuant excels in 1-bit quantization, outperforming the previous state-of-the-art method by a huge gap. We also implement efficient CUDA kernels for NSNQuant, and verify that NSNQuant achieves a 3×3$\times$3 × speedup compared to the FP16 baseline.
@@ Acknowledgements
We are grateful to VESSL AI for providing the compute resources used for the experiments in this work.
@@ A.1 Proof of Lemma 1
4. Hanson–Wright Since $s$ is a Rademacher vector, it is a sub-gaussian vector. Applying the Hanson-Wright inequality [34], for any $u>0$
@@
5. Tail parameter. Choose $t=\Gamma\beta_{\alpha}$, $\beta_{\alpha}=\frac{1}{d}\sqrt{\ln(2/\alpha)/c}$. Exponent becomes $-\ln(2/\alpha)$; hence $\Pr(|f(h)|>t)\leq\alpha$.
@@
[…]
@@ A.2 Off-diagonal Frobenius norms of covariance
Figure 7: The min-max range of off-diagonal covariances of keys (top) and values (bottom) in the first layer of LLaMA2-7B. Both key and value suffer from outliers in covariance matrices. We use the first sample from the WikiText-2 for visualization.
@@ C.1 Effects of using randomized Hadamard transform (RHT)
From Lemma 1, we find that adopting randomized Hadamard transform after the NSN process gives theoretical bounds to the variance of each channel. However, since the covariances between channels tend to have uniform signs, we find that using the naive Hadamard transform works well enough. As shown in Table 6, both transforms give similar results. Since RHT needs more parameters and computations, we use the naive Hadamard transform in the final design.
Table 6: Perplexity of LLaMA2-7B on WikiText-2 and C4 with NSNQuant-2b using different types of Hadamard transform.
Method WikiText-2 C4
Hadamard transform 5.285 6.856
RHT 5.290 6.854
@@ C.2 Effects of residual size
We fix the residual size to 64 in our main experiments, for fair evaluation across different methods. However, the residual size is an important hyperparameter that determines the number of full precision caches. Therefore, we evaluate NSNQuant with 3 different residual sizes—32, 64, 128—with LLaMA3.1-8B-Instruct. Since LongBench is not effective to show performance differences in 2-bit regime, we use GSM8K, HumanEval, and CoQA. The result is shown in Table 7. For NSNQuant-1b, we observe that performance is highly sensitive to the choice of residual size. A residual size of 128 yields the best results across most benchmarks, while a residual size of 32 performs the worst. In contrast, NSNQuant-2b shows comparable performance with different residual sizes. These results suggest that NSNQuant-2b produces higher-quality quantizations, making it more robust to variations in residual size.
We also report the memory usage and throughput with different residual sizes in Figure 8. The result shows that the methods with smaller residual sizes require less memory and achieve higher throughput. Note that the impact of residual size to memory and throughput will decrease as the sequence length gets longer.
@@ C.4 Effects of codebook tuning
As mentioned in section 3.4, we propose to fine-tune a codebook obtained from the K-Means algorithm. Since the error of scale adjustment only depends on the angle between $v$ and $v_{Q}$, we set the objective to minimize the cosine distance between them. The perplexity result with LLaMA2-7B on WikiText-2 is presented in Table 9. The result shows that fine-tuning improves quantization quality in both NSNQuant-1b and NSNQuant-2b, and matches the performance of E8P [37] in 2-bit quantization.
@@ C.5 Effects of pre-RoPE NSN
As visualized in Figure 2, NSNQuant applies NSN to keys before the RoPE. However, it seems natural to apply NSN after RoPE since RoPE may induce noises to the channel-wise mean. The PPL evaluation results in both settings are presented in Table 10. The pre-RoPE NSN achieves slightly lower PPL than the post-RoPE NSN. We attribute this to two reasons. First, the effect of RoPE on the channel-wise mean is minimal, as the tokens within the residual share the similar rotating angles. Second, computing full-precision RoPE for $o$ is beneficial, since the RoPE rotation matrix contains important position information.
@@ D.1 Replacing the Shift step with Weiszfeld algorithm
As noted in Section 3.2, the final normalization step of NSN can introduce a small bias, leaving the channel-wise mean slightly off zero. To verify that it barely affects the quantization quality, we provide a mathematical explanation and an experiment where we totally remove such bias by adopting Weiszfeld algorithm.
@@
[…]
@@
We compute $o_{\star}$ with the Weiszfeld algorithm—the standard iterative solver for geometric medians—and subtract it instead of the arithmetic mean in NSN’s second step. The resulting variant keeps the post-normalization mean essentially at zero. Table 12 presents the perplexity evaluation results on WikiText-2 and C4 with LLaMA-2-7B. The change yields only marginal negative effects despite the additional cost from the iterative updates. Hence, we keep the original mean-subtraction step in the final NSN design.
@@ D.2 Replacing the second Normalization step with channel-wise scaling
A straightforward way to standardize the output distribution is to simply shift and scale in a channel dimension. Therefore, we try replacing our third step—token-wise normalization—with channel-wise scaling, where sample standard deviation is divided from each channel. We also move the location of normalizations next to the Hadamard transform so that it does not affect channel-wise statistics. The result is shown in Table 13. Using channel-wise scaling instead of token-wise normalization is not as effective as token-wise normalization. This is because channel-wise scaling does not suppress outlier tokens, especially the attention sink token. For example, using the first sample from WikiText-2 dataset, the average norm of the first token in the first layer is 19.1, which is nearly twice as large as the overall average (11.3). Considering a ball-shaped property of our codebook, quantization error would be large for these tokens since their magnitudes are far from zero. On the other hand, our token-wise normalization effective regulates the scale, making our codebook work effectively.
@@ E.3 Additional explanation on baselines
KIVI [30] is one of the pioneering works in low-bit quantization of the KV cache of LLMs. KIVI quantizes key cache along the channel dimension and value cache along the token dimension, considering its outlier patterns. To maintain the consistency with our residual policy, we use group size of 64 for keys. Also, since smaller group sizes incur large additional bits which is critical for low-bit scenario, we use 128 for values.
@@
KVQuant [19] is another pioneering work in KV cache quantization. Similar to KIVI, KVQuant quantizes key and value cache along the channel dimension and token dimension, respectively. KVQuant employs non-uniform quantization (nuq) and dense-and-sparse quantization to improve performance further. We apply Q-Norm and 1% dense-and-sparse quantization to both KVQuant-1b and KVQuant-2b to obtain the best result.
CQ [43] is our main competitor, which shares the core idea of VQ. CQ quantizes grouped channels using learned centroids, which serve as a codebook. We evaluate the performance of CQ-4c9b for 2-bit quantization, and CQ-8c10b for 1-bit quantization.

### v2

@@ 4.2 PPL evaluation
Table 2: Perplexity (PPL) evaluation results on WikiText-2 and C4 with a context length of 4096. The results of CQ reported in the original paper are marked with $\dagger$.
@@
Table 2 presents the perplexity (PPL) of various models evaluated on WikiText-2 and C4. Although our framework uses full-precision key and value during the prefilling stage, we follow the evaluation protocol of CQ and KVQuant by quantizing them all in this experiment. Despite its difference from the generation scenario, we find PPL evaluation useful because it measures quantization quality for a sequence with a single forward pass. Furthermore, the results are aligned with those in Table 17, where we report the PPL evaluation results in the generation scenario with residuals. Therefore, we report PPL results in the main table, and use them in ablation studies.
@@ 4.3 LongBench evaluation
Table 3: Evaluation results on LongBench. The task subset is selected following KIVI (liu2024kivi). More results with different models can be found in Table 18.
Model Method Bits
Qasper
QMSum
MultiNews
TREC
TriviaQA
SAMSum
LCC
RepoBench-P
Avg.
@@ 4.4 Evaluation on more datasets
4.5 Memory usage and throughput
To evaluate the efficiency of NSNQuant in terms of both memory and runtime, we measure its memory consumption and throughput. Following previous works (liu2024kivi; kwon2023efficient), we use the synthetic data to simulate the scenario of ShareGPT (sharegpt2023) where the average input length is 161 and the average generation length is 338. The evaluation is performed on a Linux server with a single A100-80GB GPU with LLaMA2-7B.
The result is illustrated in Figure 4. While FP16 baseline suffers from OOM for large batch sizes, NSNQuant-2b and NSNQuant-1b scale efficiently to larger batch sizes, achieving $4\times$ larger batch sizes and $3\times$ speedup in throughput. Note that the memory increases are not proportional to average bits since the residual is also included in the memory usage.
4.6 Latency breakdown
Table 5: Latency (ms) breakdown of NSNQuant-2b compared with the FP16 baseline. HT and VQ refer to Hadamard transform and vector quantization, respectively. More results with different configurations can be found in Table 21
. Stage Method Total NSN HT VQ Prefill FP16 1225.90 NSNQuant-2b 1707.22 155.22 15.04 326.85 Decode (per step) FP16 50.80 NSNQuant-2b 36.49 0.34 1.88 0.66
Although NSNQuant enables the use of larger batch sizes, it introduces additional computations for the Hadamard transform, vector quantization, and NSN. To better understand their impact on efficiency, we measure the latency of each operation during the forward pass. We set the batch size to 32 and the input length to 512, and generate 64 tokens to match the residual size. The latency is measured with LLaMA2-7B on a Linux server with A100-80GB, and the reported values are averaged over 100 runs.
The result is given in Table 5. Due to the additional overhead, NSNQuant has higher latency in the prefilling stage. On the other hand, in the decoding stage, it achieves lower latency since NSNQuant alleviates the memory-bound nature of the attention computation by compressing the KV cache into lower bits. This suggests that NSNQuant is particularly advantageous for decode-heavy tasks such as reasoning or code generation.
4.7 Ablation study
@@
Figure 5 shows the cosine similarity between the original vector $v$ and the quantized vector $v_{Q}:=\mathbb{C}[\operatorname{VQ}(v)]$ with different codebooks. The fine-tuned version gives higher cosine similarities compared to the K-Means codebook. Furthermore, the cosine similarities are consistently high in the early layers, although the standardization of NSN does not hold in these layers. Notably, the cosine similarities measured using the key-value tokens (lines with markers) are very similar to those measured using synthetic standard normal data (dotted lines). This suggests that NSN successfully aligns the output distribution with the standard normal distribution, and the codebook trained only on synthetic data effectively quantizes such outcomes. As a result, codebook tuning improves perplexity from 5.294 to 5.285 with NSNQuant-2b and 6.910 to 6.703 with NSNQuant-1b.
5 Related work
Recent efforts to reduce the memory and latency bottlenecks of LLMs have largely focused on weight quantization and KV cache compression. Weight quantization methods such as GPTQ (frantar2022gptq) and AWQ (lin2024awq) significantly reduce the memory footprint of model weights while preserving accuracy. In particular, vector quantization (VQ) methods like QuIP#(pmlr-v235-tseng24a), AQLM(egiazarian2024extreme), and VPTQ (liu2024vptq) enable extremely low-bit quantization of model weights. However, these approaches do not mitigate the growing memory consumption of the KV cache, which scales linearly with sequence length during long-context inference.
To address this, recent studies have explored compressing the KV cache. KVQuant (hooper2024kvquant) and KIVI (liu2024kivi) observe the presence of outlier patterns in KV representations and propose to apply channel-wise quantization for key and token-wise quantization for value. CQ (zhang2024kv) extends non-uniform quantization (nuq) to multiple channels, achieving effective 1-bit KV cache compression. Orthogonal approaches, such as H2O (zhang2023h2o) and SnapKV (li2024snapkv), reduce memory through cache eviction of uninformative tokens. Hybrid strategies like ZipCache (he2024zipcache) and MiKV (yang2024no) selectively maintain high-precision cache for important tokens while applying low-bit quantization elsewhere. Our method uses the same precision for all tokens, but future work could incorporate token importance as in hybrid approaches.
@@ 6 Conclusion
In this work, we propose NSNQuant, a calibration-free vector quantization (VQ) method for compressing KV cache of LLMs. NSNQuant effectively aligns the token distribution with the standard normal distribution through the NSN (Normalize-Shift-Normalize) process, enabling the use of the specialized codebook for the standard normal distribution. Through extensive experiments, we demonstrate that unlike calibration-based VQ methods, NSNQuant generalizes well across different tasks and datasets. In particular, NSNQuant excels in 1-bit quantization, outperforming the previous state-of-the-art method by a huge gap. We also implement efficient CUDA kernels for NSNQuant, and verify that NSNQuant achieves a $3\times$ speedup compared to the FP16 baseline.
@@ Acknowledgements
This work was supported by SK Hynix, Samsung Advanced Institute of Technology (SAIT), IITP/NRF grants funded by the Korean government (MSIT, 2021-0-00105 Development of Model Compression Framework for Scalable On-Device AI Computing on Edge Applications, 2021-0-01343 Artificial Intelligence Graduate School Program (Seoul National University), 2021-0-02068 Artificial Intelligence Innovation Hub), and Inter-university Semiconductor Research Center (ISRC), SNU. We are also grateful to VESSL AI for providing the compute resources used for the experiments in this work.
@@ A.1 Proof of Lemma 1
4. Hanson–Wright Since $s$ is a Rademacher vector, it is a sub-gaussian vector. Applying the Hanson-Wright inequality [rudelson2013hanson], for any $u>0$
@@
5. Tail parameter. Choose $t=\Gamma\beta_{\alpha}$, $\beta_{\alpha}=\frac{1}{d}\sqrt{\ln(2/\alpha)/c}$. Exponent becomes $\ln(2/\alpha)$; hence $\Pr(|f(h)|>t)\leq\alpha$.
@@
$\operatorname{Var}(Y_{i})\in\bigl[\,1-\varepsilon-\Gamma\beta_{\alpha},\;1+\Gamma\beta_{\alpha}\bigr].$
@@ A.2 Off-diagonal Frobenius norms of covariance
Figure 7: The min-max range of off-diagonal covariances of keys (top) and values (bottom) in the first layer of LLaMA2-7B. For both key and value, some channels suffer from outliers in covariance matrices. We use the first sample from the WikiText-2 for visualization.
@@ C.1 Effects of using randomized Hadamard transform (RHT)
From Lemma 1, we find that adopting randomized Hadamard transform (RHT) after the NSN process gives theoretical bounds to the variance of each channel. However, since the covariances between channels tend to have uniform signs, we find that using the naive Hadamard transform works well enough. As shown in Table 7, both transforms give similar results. Since RHT needs more parameters and computations, we use the naive Hadamard transform in the final design.
Table 7: Perplexity of LLaMA2-7B on WikiText-2 and C4 with NSNQuant-2b using different types of Hadamard transform. We follow the settings from the main experiments for WikiText-2 and C4. For LCC and SAMSum, we truncate each data sample to 4096 tokens if needed.
Method WikiText-2 C4 LCC SAMSum
Hadamard transform 5.285 6.856 2.128 6.466
RHT 5.290 6.854 2.128 6.461
@@ C.2 Effects of residual size
We fix the residual size to 64 in our main experiments, for fair evaluation across different methods. However, the residual size is an important hyperparameter that determines the number of full precision caches. Therefore, we evaluate NSNQuant with 3 different residual sizes (32, 64, 128) with LLaMA3.1-8B-Instruct. Since LongBench is not effective to show performance differences in 2-bit regime, we use GSM8K, HumanEval, and CoQA. The result is shown in Table 8. For NSNQuant-1b, we observe that performance is highly sensitive to the choice of residual size. A residual size of 128 yields the best results across most benchmarks, while a residual size of 32 performs the worst. In contrast, NSNQuant-2b shows comparable performance with different residual sizes. These results suggest that NSNQuant-2b produces higher-quality quantizations, making it more robust to variations in residual size.
We also report the memory usage and throughput with different residual sizes in Figure 8. The result shows that the methods with smaller residual sizes require less memory and achieve higher throughput. Note that the impact of residual size on memory and throughput will decrease as the sequence length gets longer.
@@ C.4 Effects of codebook tuning
As mentioned in section 3.4, we propose to fine-tune a codebook obtained from the K-Means algorithm. Since the error of scale adjustment only depends on the angle between $v$ and $v_{Q}$, we set the objective to minimize the cosine distance between them. The perplexity result with LLaMA2-7B on WikiText-2 is presented in Table 10. The result shows that fine-tuning improves quantization quality in both NSNQuant-1b and NSNQuant-2b, and matches the performance of E8P [pmlr-v235-tseng24a] in 2-bit quantization.
@@ C.5 Effects of pre-RoPE NSN
As visualized in Figure 2, NSNQuant applies NSN to keys before the RoPE. However, it seems natural to apply NSN after RoPE since RoPE may induce noises to the channel-wise mean. The PPL evaluation results in both settings are presented in Table 11. The pre-RoPE NSN achieves slightly lower PPL than the post-RoPE NSN. We attribute this to two reasons. First, the effect of RoPE on the channel-wise mean is minimal, as the tokens within the residual share similar rotation angles. Second, computing full-precision RoPE for $o$ is beneficial, since the RoPE rotation matrix contains important position information.
@@ C.6 Effects of double quantization
C.7 Scaling calibration set for CQ
Since CQ uses only a very small dataset of 16 samples, a natural question arises: can CQ be further improved to outperform NSNQuant by scaling its calibration set? To answer this question, we test 2 variants of CQ where the size of the calibration set is doubled to 32. For the first variant, we use 32 samples from WikiText-2 (W32), enlarging the size of the calibration set. For the other variant, we use 16 samples from WikiText-2 and another 16 samples from C4 (W16C16). This variant not only increases the size but diversifies the calibration set, covering a wider range of input distribution. We use LLaMA3-8B-Instruct model to test performance on the three downstream tasks: HumanEval, GSM8K, and CoQA. We denote two variants as W32 and W16C16, respectively.
The result is shown in Table 13. While W32 shows only marginal improvement from the baseline, W16C16 provides clearer benefits. It suggests that to improve the robustness of CQ, it is important to diversify the calibration set, rather than just scaling it up. In addition, despite improved performance, NSNQuant still shows better results. For example, on HumanEval, CQ-8c10b still suffers from significant performance degradation since its input distribution (i.e. codes) is not covered by either WikiText-2 or C4.
We believe that further scaling of the calibration set of CQ will improve its generalization ability. However, since CQ needs gradient computation, the calibration process incurs comparable cost to fine-tuning, as the calibration set grows. Moreover, it’s challenging to prepare the calibration set that adequately covers the full diversity of real-world input distributions (e.g. different languages). Therefore, we believe that NSNQuant remains a highly practical and robust option in general scenarios, especially when broad generalization is required
Table 13: Performance comparison of NSNQuant and CQ baselines on HumanEval, GSM8K (8-shot, CoT), and CoQA.
Method HumanEval GSM8K (8-shot, CoT) CoQA
FP16 28.05 76.62 61.47
NSNQuant-2b 32.93 74.53 61.62
CQ-4c9b 23.78 72.55 60.60
CQ-4c9b (W32) 25.00 72.93 60.38
CQ-4c9b (W16C16) 29.27 73.16 60.98
NSNQuant-1b 29.27 61.64 60.28
CQ-8c10b 13.41 24.26 53.38
CQ-8c10b (W32) 15.24 21.76 53.37
CQ-8c10b (W16C16) 12.20 46.47 54.07
@@ D.1 Replacing the Shift step with Weiszfeld algorithm
As noted in Section 3.2, the final normalization step of NSN can introduce a small bias, leaving the channel‑wise mean slightly off zero. To verify that it barely affects the quantization quality, we provide a mathematical explanation and an experiment where we totally remove such bias by adopting Weiszfeld algorithm.
@@
$\mathcal{L}(o):=\sum_{i=1}^{\operatorname{rs}}\frac{t_{i}-o}{\lVert t_{i}-o\rVert}\;=\;\mathbf{0}.$ (2)
@@
We compute $o_{\star}$ with the Weiszfeld algorithm—the standard iterative solver for geometric medians—and subtract it instead of the arithmetic mean in NSN’s second step. The resulting variant keeps the post‑normalization mean essentially at zero. Table 14 presents the perplexity evaluation results on WikiText‑2 and C4 with LLaMA2‑7B. The change yields only marginal negative effects despite the additional cost from the iterative updates. Hence, we keep the original mean‑subtraction step in the final NSN design.
@@ D.2 Replacing the second Normalization step with channel-wise scaling
A straightforward way to standardize the output distribution is to simply shift and scale in a channel dimension. Therefore, we try replacing our third step—token-wise normalization—with channel-wise scaling, where sample standard deviation is divided from each channel. We also move the location of normalizations next to the Hadamard transform so that it does not affect channel-wise statistics. The result is shown in Table 15. Using channel-wise scaling instead of token-wise normalization is not as effective as token-wise normalization. This is because channel-wise scaling does not suppress outlier tokens, especially the attention sink token. For example, using the first sample from WikiText-2 dataset, the average norm of the first token in the first layer is 19.1, which is nearly twice as large as the overall average (11.3). Considering a ball-shaped property of our codebook, quantization error would be large for these tokens since their magnitudes are far from zero. On the other hand, our token-wise normalization effectively regulates the scale, making our codebook work effectively.
@@ E.3 Additional explanation on baselines
KIVI [liu2024kivi] is one of the pioneering works in low-bit quantization of the KV cache of LLMs. KIVI quantizes key cache along the channel dimension and value cache along the token dimension, considering its outlier patterns. To maintain consistency with our residual policy, we use group size of 64 for keys. Also, since smaller group sizes incur large additional bits which is critical for low-bit scenario, we use 128 for values.
@@
KVQuant [hooper2024kvquant] is another pioneering work in KV cache quantization. Similar to KIVI, KVQuant quantizes key and value cache along the channel dimension and token dimension, respectively. KVQuant employs non-uniform quantization (nuq) and dense-and-sparse quantization to improve performance further. We apply Q-Norm and 1% dense-and-sparse quantization to both KVQuant-1b and KVQuant-2b to obtain the best result.
CQ [zhang2024kv] is our main competitor, which shares the core idea of VQ. CQ quantizes grouped channels using learned centroids, which serve as a codebook. We evaluate the performance of CQ-4c9b for 2-bit quantization, and CQ-8c10b for 1-bit quantization.